# Safe-CLIP: Removing NSFW Concepts from Vision-and-Language Models

### v2

@@ Abstract
Large-scale vision-and-language models, such as CLIP, are typically trained on web-scale data, which can introduce inappropriate content and lead to the development of unsafe and biased behavior. This, in turn, hampers their applicability in sensitive and trustworthy contexts and could raise significant concerns in their adoption. Our research introduces a novel approach to enhancing the safety of vision-and-language models by diminishing their sensitivity to NSFW (not safe for work) inputs. In particular, our methodology seeks to sever “toxic” linguistic and visual concepts, unlearning the linkage between unsafe linguistic or visual items and unsafe regions of the embedding space. We show how this can be done by fine-tuning a CLIP model on synthetic data obtained from a large language model trained to convert between safe and unsafe sentences, and a text-to-image generator. We conduct extensive experiments on the resulting embedding space for cross-modal retrieval, text-to-image, and image-to-text generation, where we show that our model can be remarkably employed with pre-trained generative models. Our source code and trained models are available at: https://github.com/aimagelab/safe-clip. $*{}^{*}$qual contribution.
@@ 2 Related Work
Removing concepts from vision-and-language models. Removing content from AI models has been recently gaining increasing attention, with techniques spanning from complete model retraining or fine-tuning to machine unlearning [6, 17, 18] and differential privacy [19]. Some of these attempts have been considering text-to-image models and have aimed at deleting styles, concepts, or objects [25, 54]. Recently, Schramowski et al. [42] introduced a technique to steer the generation away from NSFW areas, defined by a finite and fixed set of concepts. NSFW concepts are encoded with the input prompt at inference time and the NSFW embedding is used as negative guidance. Later, Gandikota et al. [15] proposed a fine-tuning method that can erase a visual concept given only its name and using negative guidance as a teacher.
@@ 3.1 Building the ViSU dataset
Overview. In order to modify the CLIP model to avoid the representation of inappropriate content, our methodology requires a dataset comprising quadruplets of safe and unsafe (i.e. NSFW) images and sentences, denoted as:
$\mathcal{D}=\{(v_{i},t_{i},v_{i}^{\star},t_{i}^{\star}),i=1,...,N\},$ (3)
where $v_{i}$ indicates a safe image, $t_{i}$ its corresponding sentence, and the unsafe image $v_{i}^{\star}$ and unsafe sentence $t_{i}^{\star}$ are “paired” to convey a similar semantic meaning of their safe counterparts. For instance, $t_{i}$ can be considered as the sanitized version of $t_{i}^{\star}$, expressing a similar meaning without inappropriate concepts, and the same holds for their visual counterparts. As such a dataset is not available, we build $\mathcal{D}$ with an automatic annotation procedure where 1⃝ unsafe sentences $t_{i}^{\star}$ are automatically generated starting from cleaned sentences $t_{i}$, and 2⃝ unsafe images $v_{i}^{\star}$ are generated starting from unsafe sentences $t_{i}^{\star}$.
Training a conditioned NSFW textual generator. To achieve the first goal, we fine-tune a large language model (Llama 2-Chat [47]) to generate unsafe sentences starting from safe ones. In particular, we employ a set of 100 manually-curated safe-unsafe pairs, building these as a mixture of manually written pairs and sentences generated automatically with Vicuna [8]. To ensure that the dataset provides proper supervision, we follow the definition of NSFW content of [42] as that of content belonging to the following twenty categories: hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty, and balance the samples of the training dataset across these categories to encourage the LLM to generate unsafe content with good variety.
@@
Generating the full ViSU dataset. Having a conditioned NSFW generator, we can generate NSFW texts starting from safe, visually relevant, sentences. Starting from NSFW sentences, we then generate corresponding NSFW images $v_{i}^{\star}$ using a diffusion-based model which has been trained on NSFW content [^2] Specifically, we use the stablediffusionapi/newrealityxl-global-nsfw model available on HuggingFace, which has a high probability of generating NSFW images (i.e. around 80% of images have been detected as NSFW in our experiments).. The overall dataset, which we term ViSU (Visual Safe-Unsafe), contains 165k quadruplets of safe and unsafe sentences and images generated starting from COCO Captions [27]. We follow the Karpathy’s splits [22] to organize the dataset into training, validation and test splits. To ensure that the resulting dataset is balanced across the twenty NSFW categories, we prompt the NSFW generator by asking to inject a specific, randomly chosen, category into the safe sentence. Sample quadruplets from our dataset are reported in the supplementary.
@@ 4.1 Experimental Setting
Datasets. Our experiments are mainly conducted on the collected ViSU dataset that, as previously mentioned, contains 165k quadruplets of safe and unsafe content for training. For both validation and testing we use 5k samples following the Karpathy’s COCO splits [22], randomly sampling only one safe caption among the five available for each image in the original COCO dataset. When applying our Safe-CLIP to text-to-image generative architectures, we also perform experiments on the I2P dataset [42] that is composed of 4,703 textual prompts extracted from Lexica [^3] https://lexica.art/, a collection of user-generated prompts for conditioning text-to-image diffusion models. Each prompt is associated to one of seven different categories of inappropriate content, among hate, harassment, violence, self-harm, sexual content, shocking images, illegal activity.
@@ 4.3 Evaluating the Safe-CLIP Embedding Space
Retrieval results on the ViSU test set are reported in Table 2, comparing the proposed Safe-CLIP model with the original CLIP architecture, a CLIP model trained on the DataComp dataset [13], which has undergone NSFW content cleaning, and two different baselines. Specifically, we consider a variant of our approach in which we remove the two negative cosine similarity losses (i.e. Eq. 6 and 7), and a model trained with safe data only (i.e. removing the loss functions for inappropriate content redirection reported in Eq. 5 and 6). Results are reported in terms of Recall@$k$ (R@$k$) with $k=1,10,20$, that measures the percentage of times the visual or textual item associated to the query is retrieved among the top-$k$ elements. When using unsafe sentences as queries, for each element we consider the safe image associated with the given text as the corresponding visual element. Symmetrically, when using unsafe images as queries, for each element we consider the safe text associated with the given image as the ground-truth item. Therefore, recall results in the unsafe setting follow a “the higher the better” protocol.
@@
Robustness on real NSFW images. To further analyze the safety degree of the Safe-CLIP embedding space, we perform text-to-image and image-to-text retrieval using real NSFW images as visual items. Specifically, we select inappropriate visual content from three different sources: (i) a portion of data used to train the NudeNet classifier, (ii) images crawled from the web using NSFW data source URLs [^5] https://github.com/EBazarov/nsfw_data_source_urls, and (iii) images from the Socio-Moral Image Database
@@
(SMID) [10]. While the first two sources exclusively contain nudity and pornography images, the third one includes more varied types of inappropriate images representing negative concepts such as, for example, harm, inequality, discrimination, and unfairness. Overall, we randomly sample 1,000 images from each of the NSFW data sources, selecting only those representing unsafe concepts for the SMID dataset. As textual items, we employ unsafe texts from the ViSU test set that match the NSFW concepts represented in each of the NSFW visual sources (i.e. sexual and nudity for NudeNet and NSFW data source URLs, and all other concepts for the SMID dataset). For both I2T and T2I, we employ a set of 10k randomly selected visual or textual distractors, randomly selected from the LAION-400M dataset [45].
@@ 4.4 Safe-CLIP for Text-to-Image Generation
NSFW Image LLaVA LLaVA + Safe-CLIP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.
@@ 4.5 Safe-CLIP for Image-to-Text Generation
Finally, we assess the capabilities of the Safe-CLIP visual encoder when applied to an existing multimodal LLM. We employ LLaVA [30] [^7] https://huggingface.co/liuhaotian/llava-llama-2-13b-chat-lightning-preview based on LLama 2-13B-Chat, prompted by asking to describe a given image. Results are reported in Table 5 in terms of percentage of NSFW generated texts measured with GPT-3.5 and toxicity degree computed using the Perspective API. Also for this
@@
Qualitative results. Samples of generated descriptions for NSFW images are shown in Fig. 5, comparing Safe-CLIP applied to LLaVA with the original LLaVA. Notably, Safe-CLIP prevents LLaVA from generating inappropriate text while preserving its ability to describe the visual content of the scene.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has been supported by the PNRR-M4C2 project “FAIR - Future Artificial Intelligence Research” and the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No. 101120237), both funded by the European Commission, and by the PNRR project “Italian Strengthening of Esfri RI Resilience” (ITSERR) funded by the European Union - NextGenerationEU (CUP B53C22001770006).
@@ Appendix 0.A ViSU Dataset Analysis
Table 6: Mapping of ViSU categories to I2P [42].
I2P Categories ViSU Categories
hate
hate
harassment
harassment
violence
violence, suffering, humiliation, harm, child abuse, brutality, cruelty
self-harm
suicide
sexual
sexual, nudity
shocking
bodily fluids, blood, obscene gestures
illegal activity
illegal activity, drug use, theft, vandalism, weapons
@@
Analysis on generated NSFW images. As mentioned in Sec 3.1 of the main paper, we generate NSFW images using a publicly available text-to-image generator888stablediffusionapi/newrealityxl-global-nsfw, which can generate a higher number of NSFW images compared to
Table 7: Comparison between NSFW images generated by SD v1.4 and the employed NewRealityXL model8.
@@ Appendix 0.C Additional Experimental Results
Retrieval results with CLIP ViT-L/14@336 and OpenCLIP ViT-H/14. To assess the generalization capabilities of our fine-tuning strategy, we apply it to different CLIP-based models. In particular, we employ the CLIP ViT-L/14@336 (whose visual encoder is used in the best configuration of LLaVA 1.5 and 1.6) and OpenCLIP ViT-H/14 model trained on LAION-2B [44] (whose text encoder is used in Stable Diffusion v2.0). Both models are fine-tuned with the same strategy and hyper-parameters used in the main paper. Table 9 shows the retrieval
@@
results comparing our model with the original CLIP-based models and, for OpenCLIP ViT-H/14, the two baselines described in Sec. 4.3 of the main paper. Also in this setting, Safe-CLIP demonstrates superior performance than the original model and baselines. This is further confirmed when testing the model using also real NSFW images as queries or retrievable items. The results of this experiment are reported in Table 10, replicating the experimental setting shown in Table 3 of the main paper. Overall, Safe-CLIP significantly decreases the probability of retrieving inappropriate visual content when using NSFW images from all considered NSFW visual sources as retrievable items. On the same line, Safe-CLIP can also retrieve a lower percentage of NSFW textual elements when using real NSFW images as queries compared to both considered CLIP-based backbones.
@@
Text-to-image generation results with SD v2.0. We then apply the fine-tuned version of the OpenCLIP ViT-H/14 model to Stable Diffusion v2.0999Please note that the text encoder employed in Stable Diffusion v2.0 is slightly different from the one employed in the original OpenCLIP ViT-H/14. Therefore, for this experiment, we use the fine-tuned version of the modified OpenCLIP model as used in Stable Diffusion v2.0. for the text-to-image generation task. Specifically, we replicate the experiment described in Sec. 4.4 and apply Safe-CLIP to the original Stable Diffusion v2.0 model and to other variants that either employ negative prompts or the negative guidance strategy used in SLD [42]. Results are reported in Table 11, averaging the NSFW generation probabilities over five generations with different random seeds. Overall, Safe-CLIP can contribute in almost all settings to reduce the probability of generating inappropriate images, thus further demonstrating the effectiveness of our approach. The only exception is represented by the results with SLD-Strong on the I2P dataset. We argue that the strong guidance used by this version can not be effectively combined with the fine-tuned embedding space of our Safe-CLIP model. However, it is worth noting that SLD [42] can lead to more significant degradation of the realism of generated images than Safe-CLIP (cf. Table 13).
Additional qualitative results are reported in Fig. 11 and Fig. 12, using unsafe textual prompt respectively from our ViSU dataset and the I2P benchmark. We compare images generated by the original Stable Diffusion, Stable Diffusion guided with negative prompts, SLD in its Strong variant [42], and Stable Diffusion with the proposed Safe-CLIP text encoder. While competitors often fail to generate safe images, the Stable Diffusion model augmented with Safe-CLIP not only avoids generating NSFW visual content but also is able to synthesize images that preserve the original semantic content of the input textual prompts.
Image-to-text generation results with LLaVA 1.5 and 1.6. Following the same procedure described in Sec. 4.5, we apply the safe version of CLIP ViT-L/14@336 to LLaVA 1.5 [28] and LLaVA 1.6 [29] and evaluate the probability of
@@
generating unsafe text when feeding the model with real NSFW images. Results are reported in Table 12 in terms of NSFW degree and toxicity of generated text. These results confirm the ability of Safe-CLIP to effectively reduce the inappropriateness of multimodal LLMs such as LLaVA.
Also for this setting, we report in Fig. 13 some additional qualitative results comparing the generation of the LLaVA model with and without the visual encoder of Safe-CLIP. Generated textual sentences demonstrate the effectiveness of our approach in significantly reducing the probability of generating inappropriate textual content.
@@
Specifically, to evaluate generated images, we extract 30k images and corresponding captions from the COCO validation set [27] and LAION-400M [45] and compute the Frechét Inception Distance (FID) [^10] https://github.com/GaParmar/clean-fid between real and generated image distributions and the CLIP similarity between each generated image and the corresponding textual sentence. In Table 13, we compare the results using images generated by the original Stable Diffusion v1.4 model with
@@
those generated using the text encoder of Safe-CLIP. Additionally, we include the FID score and CLIP similarity considering images generated by the SLD-Strong model [42]. As it can be noticed, using Safe-CLIP in place of the original CLIP text encoder only slightly degrades the performance on both considered datasets. Nevertheless, our solution can better preserve image quality and image-text similarity than the SLD-Strong approach, which more significantly deteriorates the performance of the original Stable Diffusion model.
To evaluate generated text, instead, we consider some evaluation benchmarks typically used to evaluate the capabilities of multimodal LLMs [5].

### v3

@@ Abstract
Large-scale vision-and-language models, such as CLIP, are typically trained on web-scale data, which can introduce inappropriate content and lead to the development of unsafe and biased behavior. This, in turn, hampers their applicability in sensitive and trustworthy contexts and could raise significant concerns in their adoption. Our research introduces a novel approach to enhancing the safety of vision-and-language models by diminishing their sensitivity to NSFW (not safe for work) inputs. In particular, our methodology seeks to sever “toxic” linguistic and visual concepts, unlearning the linkage between unsafe linguistic or visual items and unsafe regions of the embedding space. We show how this can be done by fine-tuning a CLIP model on synthetic data obtained from a large language model trained to convert between safe and unsafe sentences, and a text-to-image generator. We conduct extensive experiments on the resulting embedding space for cross-modal retrieval, text-to-image, and image-to-text generation, where we show that our model can be remarkably employed with pre-trained generative models. Our source code and trained models are available at: https://github.com/aimagelab/safe-clip. Equal contribution.
@@ 2 Related Work
Removing concepts from vision-and-language models. Removing content from AI models has been recently gaining increasing attention, with techniques spanning from complete model retraining or fine-tuning to machine unlearning [6, 17, 18, 38] and differential privacy [19]. Some of these attempts have been considering text-to-image models and have aimed at deleting styles, concepts, or objects [25, 56]. Recently, Schramowski et al. [44] introduced a technique to steer the generation away from NSFW areas, defined by a finite and fixed set of concepts. NSFW concepts are encoded with the input prompt at inference time and the NSFW embedding is used as negative guidance. Later, Gandikota et al. [15] proposed a fine-tuning method that can erase a visual concept given only its name and using negative guidance as a teacher.
@@ 3.1 Building the ViSU dataset
Overview. In order to modify CLIP to avoid the representation of inappropriate content, our methodology requires a dataset comprising quadruplets of safe and unsafe (i.e. NSFW) images and sentences, denoted as $\mathcal{D}=\{(v_{i},t_{i},v_{i}^{\star},t_{i}^{\star}),i=1,...,N\}$, where $v_{i}$ indicates a safe image, $t_{i}$ its corresponding sentence, and the unsafe image $v_{i}^{\star}$ and unsafe sentence $t_{i}^{\star}$ are “paired” to convey a similar semantic meaning of their safe counterparts. For instance, $t_{i}$ can be considered as the sanitized version of $t_{i}^{\star}$, expressing a similar meaning without inappropriate concepts, and the same holds for their visual counterparts. As such a dataset is not available, we build $\mathcal{D}$ with an automatic annotation procedure where ① unsafe sentences $t_{i}^{\star}$ are automatically generated starting from cleaned sentences $t_{i}$, and ② unsafe images $v_{i}^{\star}$ are generated starting from unsafe sentences $t_{i}^{\star}$.
Training a conditioned NSFW textual generator. To achieve the first goal, we fine-tune a large language model (Llama 2-Chat [49]) to generate unsafe sentences starting from safe ones. In particular, we employ a set of 100 manually-curated safe-unsafe pairs, building these as a mixture of manually written pairs and sentences generated automatically with Vicuna [8]. To ensure that the dataset provides proper supervision, we follow the definition of NSFW content of [44] as that of content belonging to the following twenty categories: hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, abuse, brutality, cruelty, and balance the samples of the training dataset across these categories to encourage the LLM to generate unsafe content with good variety.
@@
Generating the full ViSU dataset. Having a conditioned NSFW generator, we can generate NSFW texts starting from safe, visually relevant, sentences. Starting from NSFW sentences, we then generate corresponding NSFW images $v_{i}^{\star}$ using a diffusion-based model which has been trained on NSFW content [^2] We use the stablediffusionapi/newrealityxl-global-nsfw model available on HuggingFace, which has a high probability of generating NSFW images. . The overall dataset, which we term ViSU (Visual Safe-Unsafe), contains 165k quadruplets of safe and unsafe sentences and images generated starting from COCO Captions [27]. We follow the Karpathy’s splits [22] to organize the dataset into training, validation and test splits. To ensure that the resulting dataset is balanced across the twenty NSFW categories, we prompt the NSFW generator by asking to inject a specific, randomly chosen, category into the safe sentence. Sample quadruplets from our dataset are reported in the supplementary.
@@ 4.1 Experimental Setting
Datasets. Our experiments are mainly conducted on the collected ViSU dataset that, as previously mentioned, contains 165k quadruplets of safe and unsafe content for training. For both validation and testing we use 5k samples following the Karpathy’s COCO splits [22], randomly sampling only one safe caption among the five available for each image in the original COCO dataset. When applying our Safe-CLIP to text-to-image generative architectures, we also perform experiments on the I2P dataset [44] that is composed of 4,703 textual prompts extracted from Lexica, a collection of user-generated prompts for conditioning text-to-image diffusion models. Each prompt is associated to one of seven different categories of inappropriate content, among hate, harassment, violence, self-harm, sexual content, shocking images, illegal activity.
@@ 4.3 Evaluating the Safe-CLIP Embedding Space
Retrieval results on the ViSU test set are reported in Table 2, comparing the proposed Safe-CLIP model with the original CLIP architecture, a CLIP model trained on the DataComp dataset [13], which has undergone NSFW content cleaning, and two different baselines. Specifically, we consider a variant of our approach in which we remove the two negative cosine similarity losses (i.e. Eq. 5 and 6), and a model trained with safe data only (i.e. removing the loss functions for inappropriate content redirection reported in Eq. 4 and 5). Results are reported in terms of Recall@$k$ (R@$k$) with $k=1,10,20$, that measures the percentage of times the visual or textual item associated to the query is retrieved among the top-$k$ elements. When using unsafe sentences as queries, for each element we consider the safe image associated with the given text as the corresponding visual element. Symmetrically, when using unsafe images as queries, for each element we consider the safe text associated with the given image as the ground-truth item. Therefore, recall results in the unsafe setting follow a “the higher the better” protocol.
@@
Robustness on real NSFW images. To further analyze the safety degree of the Safe-CLIP embedding space, we perform text-to-image and image-to-text retrieval using real NSFW images as visual items. Specifically, we select inappropriate visual content from three different sources: (i) a portion of data used to train the NudeNet classifier, (ii) images crawled from the web using NSFW
@@
data source URLs [^4] https://github.com/EBazarov/nsfw_data_source_urls, and (iii) images from the Socio-Moral Image Database (SMID) [10]. While the first two sources exclusively contain nudity and pornography images, the third one includes more varied types of inappropriate images representing negative concepts such as, for example, harm, inequality, discrimination, and unfairness. Overall, we randomly sample 1,000 images from each of the NSFW data sources, selecting only those representing unsafe concepts for the SMID dataset. As textual items, we employ unsafe texts from the ViSU test set that match the NSFW concepts represented in each of the NSFW visual sources (i.e. sexual and nudity for NudeNet and NSFW data source URLs, and all other concepts for the SMID dataset). For both I2T and T2I, we employ a set of 10k randomly selected visual or textual distractors, randomly selected from the LAION-400M dataset [47].
@@ 4.5 Safe-CLIP for Image-to-Text Generation
Finally, we assess the capabilities of the Safe-CLIP visual encoder when applied to an existing multimodal LLM [5]. We employ LLaVA [30] based on LLama 2-13B-Chat, prompted by asking to describe a given image. Results are reported in Table 5 in terms of percentage of NSFW generated texts measured with GPT-3.5 and toxicity degree computed using the Perspective API. Also for this
@@ 5 Conclusion
Mitigating potential misuse of the ViSU dataset. To mitigate potential misuse of the dataset, we release it via a request form, with only the textual portion available [^6] https://huggingface.co/datasets/aimagelab/ViSU-Text. Access is granted exclusively to verified researchers, who must declare their intention to use the data solely for research purposes, explicitly committing to non-malicious use. The NSFW visual part of the dataset, due to its explicit nature, is withheld to avoid potential misuse. Nonetheless, the full reproducibility of the dataset is ensured given that we employed publicly available diffusion models and we release the generation seeds and instructions.
@@ Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has been supported by the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No. 101120237), and by the the PNRR projects “FAIR - Future Artificial Intelligence Research” (M4C2 - PE00000013) and “ITSERR - Italian Strengthening of Esfri RI Resilience” (CUP B53C22001770006), both funded by the EU - NextGenerationEU.
@@
Table 6: Mapping of ViSU categories to I2P [44].
I2P Categories ViSU Categories
hate hate
harassment harassment
violence violence, suffering, humiliation, harm, abuse, brutality, cruelty
self-harm suicide
sexual sexual, nudity
shocking bodily fluids, blood, obscene gestures
illegal activity illegal activity, drug use, theft, vandalism, weapons
@@ Appendix 0.A ViSU Dataset Analysis
Analysis on generated NSFW images. As mentioned in Sec 3.1 of the main paper, we generate NSFW images using a publicly available text-to-image generator777stablediffusionapi/newrealityxl-global-nsfw, which can generate a higher number of NSFW images compared to
Table 7: Comparison between NSFW images generated by SD v1.4 and the employed NewRealityXL model7.
@@
User study to evaluate generated data in ViSU dataset. We perform a user study on 1,000 quadruplets from ViSU, involving 20 participants. For each item, we ask to evaluate if the generated text/image is unsafe and the semantic coherence between safe-unsafe texts and unsafe text-image pairs. Results are shown in Table 8, confirming that a large portion of generated data is NSFW and that the original semantics are preserved. Also, we evaluate the automatic ranking strategy used for the DPO phase. For this experiment, we use 1,000 sentence pairs and ask the user to indicate the preferred text according to its semantic coherence with the original safe caption and its NSFW degree. Human ratings agree with the raking strategy 76.4% of the time, confirming the effectiveness of our fine-tuning strategy.
Table 8: User study results evaluating generated data on the ViSU dataset and preferences in the alignment dataset for the DPO model training.
Unsafeness Unsafeness Sem. Coherence Sem. Coherence DPO Ranking
$(\mathbf{T}^{\star})$ $(\mathbf{V}^{\star})$ $(\mathbf{T},\mathbf{T}^{\star})$ $(\mathbf{T}^{\star},\mathbf{V}^{\star})$ Accuracy
89.6% 73.2% 82.8% 82.4% 76.4%
@@ Appendix 0.C Additional Experimental Results
Retrieval results with CLIP ViT-L/14@336 and OpenCLIP ViT-H/14. To assess the generalization capabilities of our fine-tuning strategy, we apply it to different CLIP-based models. In particular, we employ the CLIP ViT-L/14@336 (whose visual encoder is used in the best configuration of LLaVA 1.5 and 1.6) and OpenCLIP ViT-H/14 model trained on LAION-2B [46] (whose text encoder is used in Stable Diffusion v2.0). Both models are fine-tuned with the same strategy and hyper-parameters used in the main paper. Table 10 shows the retrieval results comparing our model with the original CLIP-based models and, for OpenCLIP ViT-H/14, the two baselines described in Sec. 4.3 of the main paper. Also in this setting, Safe-CLIP demonstrates superior performance
@@
than the original model and baselines. This is further confirmed when testing the model using also real NSFW images as queries or retrievable items. The results of this experiment are reported in Table 11, replicating the experiment shown in Table 3 of the main paper. Overall, Safe-CLIP significantly decreases the probability of retrieving inappropriate visual content when using NSFW images from all considered NSFW sources as retrievable items. On the same line, Safe-CLIP can also retrieve a lower percentage of NSFW sentences when using real NSFW images as queries compared to both considered CLIP-based backbones.
@@
Text-to-image generation results with SD v2.0. We then apply the fine-tuned version of the OpenCLIP ViT-H/14 model to Stable Diffusion v2.0 for the text-to-image generation task. Specifically, we replicate the experiment described in Sec. 4.4 and apply Safe-CLIP to the original Stable Diffusion v2.0 model and to other variants that either employ negative prompts or the negative guidance strategy used in SLD [44]. Results are reported in Table 12, averaging the NSFW generation probabilities over five generations with different seeds. Overall, Safe-CLIP can contribute in almost all settings to reduce the probability of generating unsafe images, thus further demonstrating the effectiveness of our approach. The only exception is represented by the results with SLD-Strong on the I2P dataset. We argue that the strong guidance used by this version can not be effectively combined with the fine-tuned embedding space of our Safe-CLIP model. However, it is worth noting that SLD [44] can lead to more significant degradation of the realism of generated images than Safe-CLIP (cf. Table 14).
Additional qualitative results are reported in Fig. 10 and Fig. 11, using unsafe textual prompts respectively from our ViSU dataset and the I2P benchmark. We compare images generated by the original Stable Diffusion, Stable Diffusion guided with negative prompts, SLD in its Strong variant [44], and Stable Diffusion with the proposed Safe-CLIP text encoder. While competitors often fail to generate safe images, the Stable Diffusion model augmented with Safe-CLIP not only avoids generating NSFW visual content but also is able to synthesize images that preserve the original semantic content of the input textual prompts.
Image-to-text generation results with LLaVA 1.5 and 1.6. Following the same procedure described in Sec. 4.5, we apply the safe version of CLIP ViT-L/14@336 to LLaVA 1.5 [28] and LLaVA 1.6 [29] and evaluate the probability of generating unsafe text when feeding the model with real NSFW images. Results are reported in Table 13 in terms of NSFW degree and toxicity of generated text. These results confirm the ability of Safe-CLIP to effectively reduce the inappropriateness of multimodal LLMs such as LLaVA. Also for this setting, we report in Fig. 12 some qualitative results comparing the generation of the LLaVA model with and without the visual encoder of Safe-CLIP. Generated textual sentences demonstrate the effectiveness of our approach in significantly reducing the probability of generating inappropriate textual content.
@@
To evaluate generated images, we extract 30k images and corresponding captions from the COCO validation set [27] and LAION-400M [47] and compute the FID score [37] between real and generated image distributions and the CLIP similarity between each generated image and the corresponding textual sentence. In Table 14, we compare the results using images generated by the original Stable Diffusion v1.4 model with those generated using the text encoder of Safe-CLIP. Additionally, we include the FID score and CLIP similarity considering images generated by the SLD-Strong model [44].
@@
Notably, using Safe-CLIP in place of the original CLIP text encoder only slightly degrades the performance on both datasets. Nevertheless, our solution can better preserve image quality and image-text similarity than the SLD-Strong approach, which more significantly deteriorates the performance of the original Stable Diffusion model.
To evaluate generated text, instead, we consider some evaluation benchmarks typically used to evaluate the capabilities of multimodal LLMs.